OpenStudy (anonymous):

A couch is 94 inches long. What is the length in feet and inches? (1 medal) 7 ft 10 in. 8 ft 2 in. 9 ft 2 in. 12 ft 10 in.

OpenStudy (anonymous):

12 inches=1 feet.. so 1 inch=1/12 feet so you can get 94 inches=(1/12)*94 feet here the remainder will be inches and the quotient will be feet

OpenStudy (anonymous):

I don't get it... ;-;

OpenStudy (notamathgenius):

Divide 94 into 12 

OpenStudy (anonymous):

And then.. *I got 7.83*

OpenStudy (notamathgenius):

Hmm does it want to the nearest inch?

OpenStudy (anonymous):

No, has to be exact length. Feet and inches.

OpenStudy (notamathgenius):

Well something is wrong then. If its exact its 7.83333333 And 7ft 10in is the closet you can get to that.

OpenStudy (anonymous):

OK, thanks.

OpenStudy (notamathgenius):

Np

OpenStudy (anonymous):

12*7=84 94-84=10 which is the remainder so 7 feet 10 inches

OpenStudy (notamathgenius):

I knew something was wrong.